The measure of the area of a rectangle is x2 -19x+48. If the dimensions of the rectangle are represented polynomials, find the demensions oftherectangle and then find its perimeter.
Answer by ankor@dixie-net.com(22740)

The measure of the area of a rectangle is x^2 -19x+48. If the dimensions of the
rectangle are represented polynomials, find the dimensions and then find it's perimeter.
:
Factor the equation to get the length and width
x^2 - 19x + 48 = 0
(x-16)(x-3) = 0
x = 16, the length
x = 3, the width
:
P = 2(16) + 2(3)
P = 32 + 6
P = 38 units is the perimeter